OpenStudy (anonymous):

the tenth eleventh and twelfth terms of a sequence are shown in the table below: term number:10|11|12| term: 21|24|27 which of the following shows the first five terms of sequence?

OpenStudy (anonymous):

a.)9,6,3,0,-3 b.)-6,-3,0,3,6 c.)6,3,0,-3,-5 d.)-9,-6,-3,0,3

OpenStudy (wolfe8):

Can you read this and do it? http://www.mathsisfun.com/algebra/sequences-sums-arithmetic.html You will set up at least 2 equations and find a. Then you can find the 2nd to 5th terms. Or you can do this manually and subtract the common difference backwards.

OpenStudy (campbell_st):

well look at the 10th term, the sequence is arithmetic with a common difference d = 3 the formula for a term in an arithmetic sequence is \[a_{n} =a +(n - 1)d\] a = 1st term, n = number of terms and d = common difference. so you can say \[21 = a +(10 - 1) \times 3\] solve for a, the 1st term. when you know a, the 1st term just add 3 for the 2nd and then add another 3 for the 3rd term hope it helps

OpenStudy (wolfe8):

@campbell_st Not sure if this user can see the equation notations but I cannot.

OpenStudy (campbell_st):

and no need for wolfram alpha

OpenStudy (anonymous):

im still confused

OpenStudy (wolfe8):

Manually, you can do a table starting from 10, 11, 12 like this: http://sketchtoy.com/58445088 and so on. It's a desperate move.

OpenStudy (wolfe8):

I love how my lines were straight in the beginning. (Free hand btw)

OpenStudy (anonymous):

so its b?

OpenStudy (wolfe8):

Yep looks like it.

OpenStudy (campbell_st):

ok... so do you think you can solve 21 = a + 27 for a... that will be the 1st term in the sequence

OpenStudy (anonymous):

first term is negative 6

OpenStudy (campbell_st):

correct... and you are adding 3 to get the next term... then keep adding 3... that as simple as it is

OpenStudy (anonymous):

thanks guys :)

OpenStudy (wolfe8):

You're welcome. Have a good night.